OpenStudy (anonymous):

Convert to a heavyside function: h(t): 1, pi==2pi I'm still have some problems trying to convert a step function into a heavyside function. I can do the LHS laplace transform but I can't get any further unless I convert g(t)

OpenStudy (lgbasallote):

how does this look like? \[\huge f(x) = \cases{1 \qquad \pi \le t \le 2\pi \\\\ 0 < \pi \quad \text & \quad t \ge 2\pi}\] ??

OpenStudy (lgbasallote):

\[\huge f(x) = \cases{1 \qquad \pi \le t \le 2\pi \\\\ 0 \qquad 0 \le t< \pi \quad \text & \quad t \ge 2\pi}\] is that how your question looks like?

OpenStudy (lgbasallote):

that second part doesn't seem to make sense...

OpenStudy (anonymous):

yeah, except for 1 pi=<t < 2pi

OpenStudy (lgbasallote):

can you click that draw button and draw how the question actually looks like?

OpenStudy (anonymous):

I'll take a picture, I can't seem to get the draw function to work properly for me.

OpenStudy (anonymous):

It's the one in the center.

OpenStudy (anonymous):

something like H(t-pi) - H(t-2pi) you mean?

OpenStudy (anonymous):

yeah, I don't really understand how to convert a step function into that format.

OpenStudy (anonymous):

Heaviside steps up to one when t=0, shift it to t=pi (that's the first term) Now make it drop back to zero: subtract a Heaviside that's been time shifted to 2pi

OpenStudy (anonymous):

hmm I see... I kind of understand now.

OpenStudy (anonymous):

Paul's Notes goes over it... it's pretty simple really:) http://tutorial.math.lamar.edu/Classes/DE/StepFunctions.aspx is this for diff. eq.? or an engineering class?

OpenStudy (anonymous):

It's for a differentials class :\

OpenStudy (anonymous):

opps forgot to close this thread. I finally understand the heaviside format but instead of h(t-c) I used the u_c(t) format -> e^cs/s.